A 43.29 gram sample of cobalt is heated in the presence of excess oxygen. A metal oxide is formed with a mass of 60.92 g. Determine the empirical formula of the metal oxide.
Answer #1

we have mass of each elements as:

Co: 43.29 g

O: 17.63 g

Divide by molar mass to get number of moles of each:

Co: 43.29/58.93 = 0.7346

O: 17.63/16.0 = 1.1019

Divide by smallest:

Co: 0.7346/0.7346 = 1

O: 1.1019/0.7346 = 1.5

multiply by 2 to get simplest whole number ratio:

Co: 1*2 = 2

O: 1.5*2 = 3

So empirical formula is:Co2O3

Answer: Co2O3